• End of section 4.3:

“The GBA represents the partitioning graph […] and the processors interconnection graph into a simpler space”.

Not grammatical.

• Section 5, para 2:

“[…] approximated by a linear system with 11,676 number of equations”

(my italics). Not grammatical.

• Section 5, para 3, sentence 3: “Also” misspelt.

• Section 5.1, para 4, sentence 2: “and” misspelt.

• Section 5.1, para 5, sentence 5:

“[…] the total sum […]”. Redundant use of “total”.